A point charge has a charge of 9.00 x 10-12 C. At what distance from the point charge is the electrical potential (a) 13.0 V? (b.) 25.0 v?



1
Expert's answer
2022-03-31T12:26:51-0400

(a)We know that electric potential

"V=\frac{kq}{r}\\r=\frac{kq}{V}\\r=\frac{9\times10^9\times9\times10^{-12}}{13}=0.00623m"

Part (b)

"V=\frac{kq}{r}\\r=\frac{kq}{V}\\r=\frac{9\times10^9\times9\times10^{-12}}{25}=0.00324m"